A child wanders slowly down a circular staircase from the top of a tower. With x,y,z in feet and the origin at the base of the tower, her position t minutes from the start is given by
x=15cost,y=15sint,z=120−5t.
t.
(a) How tall is the tower?
height =
ft
(b) When does the child reach the bottom?
time =
minutes
(c) What is her speed at time tt?
speed =
ft/min
(d) What is her acceleration at time tt?
acceleration =
ft/min2

Answer by jim_thompson5910(35256)

(a)

Plug t = 0 into the z equation to get

z = 120 - 5*t
z = 120 - 5*0
z = 120 - 0
z = 120

This means that when the time is t = 0 minutes, they are at z = 120.

The tower is 120 feet tall.

--------------------------------------------------------------------------------------------------------
(b)

The child will reach the bottom when z = 0

Plug z = 0 into the z equation and solve for t

z = 120 - 5*t
0 = 120 - 5*t
0 + 5*t = 120 - 5*t + 5*t
5*t = 120
5*t/5 = 120/5
t = 24

The child will reach the bottom at t = 24 minutes

--------------------------------------------------------------------------------------------------------
(c)

Unfortunately I don't have enough information to answer this problem. You wrote "tt" but didn't write any values after that. Please update the problem. Thank you.

--------------------------------------------------------------------------------------------------------
(d)

As in part (c), the same issue happens with this part as well. There isn't enough information. Please update.